OpenStudy (anonymous):

An architect needs to determine the slope between two points on a ski lift. The two points have been identified as (15, 45) and (150, 60), where x is the horizontal distance and y is the vertical distance from the bottom of the lift. Assuming the lift runs in a straight line, what is the slope of the line between the two points? (Write your answer in simplest form, using / for a fraction bar if needed

OpenStudy (btaylor):

\[slope = \frac{rise}{run} = \frac{\Delta y}{\Delta x}=\frac{y_2-y_1}{x_2-x_1} = \frac{60-45}{150-15} = \frac{15}{135}=\frac{1}{9}\]